If a particle moving with velocity −4 m/s i^ and acceleration is 2 m/s2i^. What is distance travelled by the paricle in 6 sec.

Grade:12

1 Answers

Arun
6 years ago
Dear student
a = -2
u = 4
Now
From v = u +a*t
0 = 4 - 2t
t = 2sec
 
Hence after 2 sec it will stop as speed will be zero.
Now
We have to find the distance covered in 2sec.
S = u*t + (1/2) *a *t²
S = 8 - (1/2)* 2 *2²
S = 8 -4 = 4 m
 
Regards
Arun (askIITians forum expert)